solve the following systems by using either addition or subtraction. if a unique solution does not exist, state whether the system is dependent or inconsistent.
10x+2y=7
y=-5x+3

Answer by uma(370)

10x+2y=7-------------(1)
y=-5x+3--------------(2)
(2) can be written as 5x+y =3
Multiplying this by 2 we get,
10x+2y=6
10x+2y=7---------(1)
On subtracting these equations we find that, both the variables get cancelled.
==> the given system does not have a solution.
In other words the lines represented by these equations are parallel.